OpenStudy (buttermequeasy):

I f f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)] −13 −25 25 17

OpenStudy (anonymous):

evaluate g(-2) and put that into f(x)

OpenStudy (anonymous):

so essentially do g(-2) to get a scalar correct? that will be your your x in f(x)

OpenStudy (anonymous):

tell me whatcha get

OpenStudy (anonymous):

or not... answer is 25 lol

OpenStudy (anonymous):

g(-2) = -13 f(-13) = 26-1 = 25

OpenStudy (anonymous):

I try to make it a collaborative experience, but some people just don't work with me =/

OpenStudy (anonymous):

aww sorry

OpenStudy (unklerhaukus):

f( g ( -2 )) = f ( 6(-2) -1 ) = -2( 6(-2) -1) -1 = 25

OpenStudy (buttermequeasy):

Sorry I was away from the computer. :|

OpenStudy (buttermequeasy):

:/